OpenStudy (brittanydosey):

17. Your fixed expenses are $1,328.90/month and you saved 4 months' worth in an emergency fund. You place half in a 45-day CD at a 4.6% APR and the remainder in a regular savings account at a 3.7% APR. How much total interest do you earn in 45 days?

OpenStudy (brittanydosey):

First, figure out how much your saved. Then figure out how much is going into the two accounts. Use the I = Prt formula for each account. Add the two interest amounts together for the final answer.

OpenStudy (brittanydosey):

@tylermcmullen23

OpenStudy (anonymous):

Is this the question you need help on?

OpenStudy (brittanydosey):

yes

OpenStudy (anonymous):

Don't know why I even asked that question since you wouldn't of posted it otherwise...

OpenStudy (anonymous):

So we are looking at an interest problem here

OpenStudy (brittanydosey):

yes

OpenStudy (anonymous):

So in the equation what is the variable "l"

OpenStudy (brittanydosey):

@hick4life

OpenStudy (brittanydosey):

that is the thing, i have no idea. .-.

OpenStudy (hick4life):

Oh ok i see where your problem is brittney

OpenStudy (anonymous):

I = 1328.9 × 0.083 × 0.12 = 13.24

OpenStudy (anonymous):

So you add up the interest together and get 8.3%

OpenStudy (anonymous):

So 13.24 is assuming its only one months worth of pay

OpenStudy (brittanydosey):

@_fluffeh_

OpenStudy (brittanydosey):

so then is it 13.24?

OpenStudy (anonymous):

Hold on I'm going about this wrong. I forgot that it's two separate funds

OpenStudy (anonymous):

I = 2657.8 × 0.037 × 0.12 = $ 11.80 + I = 2657.8 × 0.047 × 0.12 = $14.99

OpenStudy (anonymous):

or 11.8 + 14.99

OpenStudy (brittanydosey):

26.79

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

it says that you have to figure out how much there is after 4 month and put half of that into the 45 day CD thing and the other half into the other account.. @brittanydosey was 26.79 correct?

OpenStudy (brittanydosey):

yes

OpenStudy (anonymous):

Medal?

OpenStudy (brittanydosey):

i gave you one

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

Sweet thanks guys @brittanydosey @tylermcmullen23